What are the coordinates of the suspect's car at its final destination?

(xy) = 

(b)

What is the distance (in mi) traveled by the suspect?


(c)

What is the distance (in mi) as the crow flies between the original position and the final position of the suspect's car? (Round your answer to two decimal places.)

 


1
Expert's answer
2021-09-02T07:25:39-0400

(a)  5 mi due east: "(0,0)\to(5,0)"

4 mi due north: "(5,0)\to(5,4)"

1 mi due west: "(5,4)\to(4,4)"


The coordinates of the suspect's car at its final destination are "(4,4)"


(b)


"5\ mi+4\ mi+1\ mi=10\ mi"

The distance traveled by the suspect is "10\ mi."


(c)


"d=\sqrt{(4-0)^2+(4-0)^2}=4\sqrt{2}\approx5.66(mi)"

The distance as the crow flies between the original position and the final position of the suspect's car is "5.66" mi.